OpenStudy (anonymous):

simplify sqrt 150x^5y^8

OpenStudy (angela210793):

5x^2y^8sqr(6x)

OpenStudy (anonymous):

how did you get that

OpenStudy (radar):

\[\sqrt{150x ^{5}y ^{8}}\]

OpenStudy (angela210793):

sqr(25*6*x^4*x*[y^(4)]^2)=5x^2y^4sqr(6x) sorry

OpenStudy (anonymous):

what does that say can you make it a formula so i can know what you are talking about please

OpenStudy (angela210793):

i have expressed 150=25*6 x5=x^4*x y^8=(y^4)^2

OpenStudy (anonymous):

ok and how would i get the answer based on what you have given me

OpenStudy (angela210793):

by sqrooting them sqr(25*6)=5sqr6 sqr(x^4*x)=x^2sqrx sqr{[y^4]^2}=y^4 W8 for Radar's answer though

OpenStudy (angela210793):

whr did sqr30x come from?

OpenStudy (angela210793):

shouldn't it b sqt6x?

OpenStudy (radar):

Those had to remain within the radical as they were not perfect squares.

OpenStudy (angela210793):

only 6 remains under sqr....right? O.o

OpenStudy (radar):

You are right, I mistook the 5 for a constant in the radical rather than an exponent, although I treated it as an exponent also lol Duh!\[5x ^{2}y ^{4}\sqrt{6x}\]

OpenStudy (angela210793):

phew :D :D

OpenStudy (radar):

Since it was \[x ^{5}=x ^{4}x ^{1}\]You still had a x left in the radical;

OpenStudy (anonymous):

so it is \[5x^2y^4\sqrt{6x}\]

OpenStudy (anonymous):

that is the correct answer

OpenStudy (angela210793):

yes tht's it :)

OpenStudy (radar):

I agree also, good luck with these things.

OpenStudy (anonymous):

thanks